A ship travels from New York to Southampton, England at 33 miles per hour. A plane traveling 605 miles per hour takes 104 hours less time to make the same trip. How far apart are the two cities?
Answer by Paul(988)

PLane's speed = 605mph
Ship's speed = 33mph

Since it takes 104 less ONLY for the plane. its -->x-104
Since its for plane so:
605(x-104)

And ship times the original times. --->33x
Equation:

605(x-104)=33x
605x-62920=33x
-572x=-62920
x=110

605(110-104)=33(110)
3630=3630

Hence, the cities are 3630 miles apart.
Paul.